OpenStudy (anonymous):

After checking the assumptions, someone has computed the following 93% confidence interval for a set of sample data: 38< μ<57. What is the margin of error for this interval? What is the sample mean?

OpenStudy (anonymous):

im not sure how to go about this?

OpenStudy (kropot72):

The equation to use is as follows: \[\bar{x}-1.81\frac{\sigma}{\sqrt{n}}<\mu < \bar{x}+1.81\frac{\sigma}{\sqrt{n}}\] You are given the values of the expressions on either side of < mu < so you can set up a pair of equations to solve for the value of \[\frac{\sigma}{\sqrt{n}}\] The margin of error is found from calculating to value of \[1.81\times \frac{\sigma}{\sqrt{n}}\]

OpenStudy (anonymous):

all the examples in my book give me the n and standard deviation how am i to find those values?

OpenStudy (anonymous):

based on the limited information, I believe the mean of the sample is the middle of 38 and 57: 47.5 And, the margin of error is the difference between the sample mean (47.5) and the top of the interval (57): 9.5

OpenStudy (kropot72):

\[\bar{x}+1.81\frac{\sigma}{\sqrt{n}}=57\ ..........(1)\] \[\bar{x}-1.81\frac{\sigma}{\sqrt{n}}=38\ ...........(2)\] Subtracting equation (2) from equation (1) gives \[3.62\frac{\sigma}{\sqrt{n}}=19\] Therefore \[\frac{\sigma}{\sqrt{n}}=\frac{19}{3.62}=5.2486\] and the margin of error is 5.2486 * 1.81 = 9.5 The sample mean is found by adding equations (1) and (2) giving \[2\times \bar{x}=95\] \[\bar{x}=\frac{95}{2}=47.5\]